OpenStudy (anonymous):

Which equation goes with the t-chart below? x y -2 13 -1 7 0 5 1 7 2 13 y= 5x2- 2 y= 5x2+ 2 y= 2x2+ 5 y= 2x2- 5

OpenStudy (hba):

sorry i dont know how to solve this one

OpenStudy (anonymous):

lol its ok

OpenStudy (hba):

:)

ganeshie8 (ganeshie8):

see the third row in t-chart

ganeshie8 (ganeshie8):

0 5

ganeshie8 (ganeshie8):

it says, when x =0, y = 5 right ?

ganeshie8 (ganeshie8):

put x =0 in the options, and see which option gives y = 5

ganeshie8 (ganeshie8):

can you try that..

OpenStudy (anonymous):

a ?

ganeshie8 (ganeshie8):

how, y= 5x2- 2 if you put x = 0 y = 5(0) -2 = -2

ganeshie8 (ganeshie8):

but we are looking for 5, so its not the one

OpenStudy (anonymous):

ehh ok c?

ganeshie8 (ganeshie8):

\(\huge \checkmark\)